# DN-Splatter: Depth and Normal Priors for Gaussian Splatting and Meshing

### v2

@@ 4.3 Optimization
$\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{\text{depth}}\mathcal{L}_{\hat{D}}+\mathcal{L}_{\text{scale}}+(\underbrace{\lambda_{\text{normal}}\mathcal{L}_{\it{\hat{N}}}+\lambda_{\text{smooth}}\mathcal{L}_{\hat{N}_{\text{smooth}}}}_{\mathcal{L}_{\text{normal}}})$ (11)
where $\mathcal{L}_{\text{rgb}}$ is the original photometric loss proposed in [19]. We set $\lambda_{\text{depth}}=0.2$, $\lambda_{\text{normal}}=0.1$, and $\lambda_{\text{smooth}}=0.5$ in our experiments.
@@ 4.4 Meshing
After optimizing with our depth and normal regularization using Eq. 11, we apply Poisson surface reconstruction [18] to extract a mesh. SuGaR [12] proposes to extract a coarse mesh from the scene by projecting rays from camera views and linearly interpolating intersection points based on the value of the local density (cf. Eq. 2) derived from nearby Gaussians along a ray. However, since scenes contain thousands of Gaussians, the local density along nearby points is non-smooth, resulting in noisy surfaces (cf. Table 7). In the context of depth regularization, we have already ensured that the positions of the Gaussians are well-distributed and aligned along the surface of the scene. Therefore, we directly back-project rendered depth and normal maps from training views to create an oriented point set for meshing. We show qualitative and quantitative differences between various Poisson mesh extraction methods in Table 7.
@@ 5 Experiments
In this section, we demonstrate the proposed regularization strategy on mesh extraction and novel view synthesis results using indoor datasets.
@@
Datasets. We focus on indoor datasets and consider the following: a) MuSHRoom [38]: a real-world indoor dataset containing separate training and evaluation camera trajectories; b) ScanNet++ [56]: a real-world indoor dataset with high fidelity 3D geometry and RGB data; and c) Replica [44]: smaller real-world indoor scans.
@@
Baselines. We consider a range of baseline methods for comparison a) state-of-the-art NeRF-based method Nerfacto [45]; b) its depth regularized version Depth-Nerfacto with a direct loss on ray termination distribution for depth supervision similar to DS-NeRF [6]; c) Neusfacto [57] and MonoSDF [58] for SDF-based implicit surface reconstruction; d) baseline 3DGS Splatfacto method based on Nerfstudio v1.1.3 [45]; e) SuGaR [12] – a 3DGS variant for mesh reconstruction; and f) the recent 2DGS [16] method.
@@
Evaluation metrics. We follow standard practice and use PSNR, SSIM and LPIPS metrics for color images and report common depth metrics, similar to [51, 23, 29, 42, 33, 47, 61], to analyze depth quality for datasets containing sensor or ground truth depth data. For mesh evaluation, we follow the evaluation protocol from  [38, 49] and report Accuracy (Acc.), Completion (Comp.), Chamfer-$L_{1}$ distance (C-$L_{1}$), Normal Consistency (NC), and F-scores (F1) with a threshold of 5cm.
@@
Implementation details. The proposed approach is implemented in PyTorch [35] and gsplat [55] (v1.0.0). We train all models for 30k iterations. To obtain monocular normal cues, we propagate RGB images through the pre-trained Omnidata model [7]. For Poisson reconstruction, we extract a total of 2 million points and use a depth level of 9 for all methods, unless otherwise stated. All meshes are extracted using back-projection of depth and/or normal maps besides Neusfacto and MonoSDF (the Marching cubes) and 2DGS (TSDF). More settings can be seen in the supplementary materials.
@@
ScanNet++: 8b5caf3398MuSHRoom: vr_roomNormalMeshNormalMeshNerfactoDepth-NerfactoMonoSDFSplatfactoSuGaR2DGSDN-SplatterGTRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Sensor Depth Algorithm Accuracy $\downarrow$ Completion $\downarrow$ Chamfer-$L_{1}$downarrow Normal Consistency $\uparrow$ F-score $\uparrow$
@@
Depth-Nerfacto [45] $\checkmark$ Poisson .0447 .0557 .0502 .7614 .6966
SDF MonoSDF [58] $\checkmark$ Marching-Cubes .0310 .0190 .0250 .8846 .9211
@@
DN-Splatter (Ours) $\checkmark$ Poisson .0239 .0194 .0216 .8822 .9243
Table 1: Mesh evaluation: MuSHRoom. The results are averaged over 6 scenes: ”coffee_room”, ”honka”, ”kokko”, ”sauna”, ”computer”, and ”vr_room”. Incorporating depth and normal priors to 3DGS optimization significantly improves mesh reconstruction on challenging real world indoor-datasets. The best and second best results are marked with bold and underline.
Sensor Depth Algorithm Accuracy $\downarrow$ Completion $\downarrow$ Chamfer-$L_{1}$downarrow Normal Consistency $\uparrow$ F-score $\uparrow$
@@
MonoSDF [58] $\checkmark$ Marching-Cubes .0303 .0573 .0438 .8881 .8577
@@
DN-Splatter (Ours) $\checkmark$ Poisson .0487 .0873 .0680 .8404 .6941
Table 2: Mesh evaluation: ScanNet++. The results are averaged over the ”b20a261fdf” and ”8b5caf3398” scenes of the dataset. The best and second best results are marked with bold and underline.
@@ 5.1 Mesh evaluation
We demonstrate the effectiveness of the proposed depth and normal regularization strategy on scene geometry by extracting meshes directly after optimization, without any additional refinement steps. In Table 1 and Table 2 we show how incorporating geometric cues enables competitive mesh extraction on scenes from the MuSHRoom [38] and ScanNet++ [56] datasets. We first observe that without the usage of monocular depth cues, neither NeRF- [45] nor Gaussian-based methods [16] work well for indoor reconstruction. Improving depth and normal estimation over the baseline Splatfacto [45] makes our method competitive even against the more computationally expensive SDF approaches [58, 57]. NeRF-based methods fail to consistently achieve similar results, with depth supervision sometimes hindering mesh performance. Fig. 3 provides a qualitative comparison of our mesh extraction approach with other NeRF and 3DGS baselines on the ScanNet++ dataset.
5.2 Novel view synthesis and depth estimation
@@
Sensor Depth Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta<1.25$uparrow PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Table 3: Depth estimation and novel view synthesis: MuSHRoom. The reported results are reported for two distinct evaluation datasets: left/right where left is a test set obtained by sampling uniformly every 10 frames within the training sequence and right is a test split obtained from a different camera trajectory with no overlap with the training sequence. Results are averaged over 5 scenes.
We show an extensive study on depth and normal supervision and its effect on novel view synthesis and depth metrics on challenging real-world scenes from MuSHRoom and ScanNet++ datasets. Table 3 demonstrates that incorporating sensor depth supervision into 3DGS enhances depth and RGB metrics in comparison to scenarios without geometric supervision. We qualitatively highlight the improvements in novel view synthesis and depth quality with the reduction of floaters and artefacts in Fig. 4.
@@
No Cues ($\mathcal{L}_{\text{rgb}}$[19, 45] 6.99 78.0 56.3 26.6 .870 12.9
@@
Table 4: Ablation of geometric supervision. Geometric cues significantly improve reconstruction quality. We report mesh, novel view, and training time metrics (Nvidia 4090 GPU, 30k iterations) on the ”VR room” sequence of the MuSHRoom dataset [38].
(a) iPhone GT[Uncaptioned image] (b) baseline[Uncaptioned image] (c) $\mathcal{L}_{\text{MSE}}$[Uncaptioned image] (d) $\mathcal{L}_{1}$[Uncaptioned image] (e) $\mathcal{L}_{\text{LogL1}}$[Uncaptioned image] (f) $\mathcal{L}_{\hat{D}}$ (Ours)[Uncaptioned image]
Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta<1.25$uparrow PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Splatfacto [45] .1481 .1345 .5122 .7602 23.41 .9111 .1316
Splatfacto + $\mathcal{L}_{\text{MSE}}$0551 .0180 .2104 .9647 23.84 .9140 .1246
Splatfacto + $\mathcal{L}_{\mathrm{1}}$0351 .0170 .1889 .9758 23.74 .9138 .1235
Splatfacto + $\mathcal{L}_{\mathrm{LogL1}}$0364 .0180 .1955 .9744 23.77 .9139 .1234
Splatfacto + $\mathcal{L}_{\hat{D}}$0285 .0162 .1790 .9782 23.61 .9122 .1269
Table 5: Ablation of depth losses: ScanNet++. We compare various depth supervision strategies. We observe that the proposed gradient aware $\mathcal{L_{\hat{D}}}$ regularizer obtains the best qualitative results, mitigating uncertainties at edges from the raw iPhone depth captures. Zoom in to see the details.
5.3 Ablation studies
Regularization strategy. We evaluate various design choices of the proposed method in Table 4. The normal loss $\mathcal{L}_{\hat{N}}$ (Eq. 9) helps align Gaussians along the scene geometry leading to a better scene geometry representation. The depth loss $\mathcal{L}_{\hat{D}}$ (Eq. 4) significantly improves reconstruction quality and novel-view synthesis in ambiguous, textureless regions which are common in indoor scenes. The normal smoothing prior $\mathcal{L}_{\text{smooth}}$ (Eq. 10) enhances normal-completeness for Poisson meshing with a minimal impact on other metrics. Although the smoothing prior’s effect on quantitative metrics is minimal, its importance is evident in the qualitative renders illustrated in Fig. 5.
Depth supervision and losses. We analyze the impact of the proposed depth loss on the ScanNet++ dataset, providing both qualitative and quantitative results in Table 5. We compare several common depth losses: $\mathcal{L}_{\text{MSE}}$, $\mathcal{L}_{\mathrm{1}}$, $\mathcal{L}_{\mathrm{LogL1}}$ [14] and our proposed edge-aware $\mathcal{L}_{\hat{D}}$ loss, examining their effect on depth and novel view synthesis metrics. Further comparisons of depth loss performance against ground truth lidar scans and definitions can be found in the supplementary material.
@@ 5.3 Ablation studies
Figure 5: Qualitative comparison of normal supervision. (a) We observe that using a direct normal loss $\mathcal{L}_{\hat{N}}$ results in non-smooth surface estimates whereas (b) using a normal smoothing prior $\mathcal{L}_{\text{smooth}}$ significantly improves DN-Splatter’s normal predictions according to (c) Omnidata  [7] predictions.
We observe that $\mathcal{L}_{\mathrm{1}}$ and $\mathcal{L}_{\mathrm{LogL1}}$ generally perform the best on color metrics, with the gradient-based logarithmic variant providing the smoothest qualitative reconstructions.
In addition, in Table 6 we evaluate our regularization strategy against other alternatives. We demonstrate that mesh reconstruction using current state-of-the-art monocular [3, 15] or multi-view [59] depth networks is far less accurate compared to reconstruction using low-resolution iPhone sensor depths. Monocular and multi-view depth estimates are aligned using the strategy outlined in Section 4.1 and Eq. 5, and mesh metrics are computed on the ScanNet++ dataset. We also investigate the patch-based Pearson Correlation loss [52] to examine relative depth supervision with monocular depth [15] estimates. Although the method exceeds naive depth supervision with monocular estimates, the reconstruction quality remains inferior compared to using iPhone depths.
Acc. $\downarrow$ Comp. $\downarrow$ C-$L_{1}$downarrow NC $\uparrow$ F-$\uparrow$

### v3

@@ 4.4 Optimization
$\mathcal{L}=\mathcal{L}_{\hat{C}}+\lambda_{\text{d}}\mathcal{L}_{\hat{D}}+\mathcal{L}_{\text{scale}}+(\underbrace{\lambda_{\text{n}}\mathcal{L}_{\it{\hat{N}}}+\lambda_{\text{s}}\mathcal{L}_{\text{smooth}}}_{\mathcal{L}_{\text{normal}}}),$ (12)
where $\mathcal{L}_{\hat{C}}$ is the original photometric loss proposed in [19]. We set $\lambda_{\text{d}}=0.2$, $\lambda_{\text{n}}=0.1$, and $\lambda_{\text{s}}=0.1$ in our experiments.
@@ 4.5 Meshing
After optimizing with our depth and normal regularization using Eq. 12, we apply Poisson surface reconstruction [18] to extract a mesh. With our regularization strategy, we ensure that the positions of Gaussians are well-distributed and aligned along the surface of the scene. We directly back-project rendered depth and normal maps from training views to create an oriented point set for meshing. Qualitative and quantitative differences between various Poisson methods and comparisons with the meshing approach proposed in SuGaR [11] are given in the Supplementary material.
@@ 5 Experiments
In this section, we demonstrate the proposed regularization strategy on mesh extraction and novel-view synthesis results using indoor datasets. We also give insight regarding various depth supervision strategies and initialization schemes enabled by sensor depth data.
@@
Datasets. We focus on indoor datasets and consider the following: a) MuSHRoom [37]: a real-world indoor dataset containing separate training and evaluation camera trajectories; and b) ScanNet++ [57]: a real-world indoor dataset with high fidelity 3D geometry and RGB data.
@@
Baselines. We consider a range of baselines including implicit NeRF and SDF based representations and explicit Gaussian based methods. We consider a) state-of-the-art NeRF-based method Nerfacto [45]; b) its depth regularized version Depth-Nerfacto with a direct loss on ray termination distribution for depth supervision similar to DS-NeRF [6]; c) Neusfacto [58] and MonoSDF [59] for SDF-based implicit surface reconstruction; d) baseline 3DGS Splatfacto method based on Nerfstudio v1.1.3 [45]; e) SuGaR [11] – a 3DGS variant for mesh reconstruction; and f) the recent 2DGS [16] method. In addition, for mesh reconstruction we consider g) traditional Poisson meshing utilizing back-projected sensor depth readings.
@@
Evaluation metrics. We follow standard practice and report PSNR, SSIM and LPIPS metrics for color images and common depth metrics, similar to [51, 23, 28, 42, 32, 47, 62], to analyze depth quality for datasets containing ground truth sensor data. For mesh evaluation, we follow evaluation protocols from  [37, 49] and report Accuracy (Acc.), Completion (Comp.), Chamfer-$L_{1}$ distance (C-$L_{1}$), Normal Consistency (NC), and F-scores (F1) with a threshold of 5cm.
@@
Implementation details. The proposed method is implemented in PyTorch [34] and gsplat [55] (v1.0.0). We train all models for 30k iterations. To obtain monocular normal cues, we propagate RGB images through the pre-trained Omnidata model [7]. We initialize the Gaussian scene using 1M back-projected points from training dataset sensor depths. For Poisson reconstruction, we extract a total of 2 million points and use a depth level of 9 for all methods, unless otherwise stated. All meshes are extracted using back-projection of depth and normal maps besides Neusfacto and MonoSDF (marching cubes) and 2DGS (TSDF). More settings can be seen in the supplementary material.
Gaussian initialization. For all Gaussian based baselines (Splatfacto, SuGaR, 2DGS, and DN-Splatter) we utilize 1M back-projected sensor depth points for initialization of the Gaussian scene unless otherwise stated. We compare sparse COLMAP [41, 40] SfM initialization and sensor depth initialization strategies in Table 5.
@@
ScanNet++: 8b5caf3398MuSHRoom: vr_roomNormalMeshNormalMeshNerfactoDepth-NerfactoMonoSDFSuGaR2DGSSplatfactoDN-SplatterGTRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Sensor Depth Loss Algorithm Accuracy $\downarrow$ Completion $\downarrow$ Chamfer-$L_{1}$downarrow Normal Consistency $\uparrow$ F-score $\uparrow$ Num GS
@@
[…]
Depth-Nerfacto [45] $\checkmark$ Poisson .0447 .0557 .0502 .7614 .6966 -
SDF MonoSDF [59] $\checkmark$ Marching-Cubes .0310 .0190 .0250 .8846 .9211 -
@@
DN-Splatter (Ours) $\checkmark$ Poisson .0239 .0194 .0216 .8822 .9243 1.18M
Table 1: Mesh evaluation: MuSHRoom. Adding depth and normal priors to 3DGS optimization greatly enhances mesh reconstruction on challenging real-world indoor datasets. We report the number of Gaussians in the final optimized scene for Gaussian-based models, with results averaged over 6 scenes: ’coffee_room’, ’honka’, ’kokko’, ’sauna’, ’computer’, and ’vr_room’. The best and second-best results are indicated with bold and underline.
Sensor Depth Loss Algorithm Accuracy $\downarrow$ Completion $\downarrow$ Chamfer-$L_{1}$downarrow Normal Consistency $\uparrow$ F-score $\uparrow$ Time (min)
@@
Depth-Nerfacto [45] $\checkmark$ Poisson .0731 .1647 .1189 .6848 .5018 8.1
[…]
MonoSDF [59] $\checkmark$ Marching-Cubes .0303 .0573 .0438 .8881 .8577 47.5
@@
DN-Splatter (Ours) $\checkmark$ Poisson .0940 .0395 .0667 .8316 .7658 36.9
Table 2: Mesh evaluation: ScanNet++. The results are averaged over the ’b20a261fdf’ and ’8b5caf3398’ scenes. The best and second best results are marked with bold and underline. Training time is reported using a Nvidia 4090 GPU.
@@ 5.1 Mesh evaluation
We demonstrate the effectiveness of the proposed regularization strategy on scene geometry by extracting meshes directly after optimization, without any additional refinement steps. In Table 1 and Table 2 we show how incorporating geometric cues enables competitive mesh extraction on scenes from the MuSHRoom [37] and ScanNet++ [57] datasets. On challenging indoor scenes, neither NeRF [45] nor Gaussian-based methods [16, 11] perform well for indoor reconstruction. Improving depth and normal estimation over the baseline Splatfacto [45] makes our method competitive even against the more computationally expensive SDF approaches [59, 58]. NeRF-based methods fail to consistently achieve similar results, with depth supervision sometimes hindering mesh performance. Fig. 3 provides a qualitative comparison of our mesh extraction approach with other baselines on the MuSHRoom and ScanNet++ datasets.
5.2 Novel-view synthesis and depth estimation
@@
Sensor Depth Loss Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta<1.25$uparrow PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Table 3: Depth estimation and novel view synthesis: MuSHRoom. The reported results are reported for two distinct evaluation datasets: left/right where left is a test set obtained by sampling uniformly every 10 frames within the training sequence and right is a test split obtained from a different camera trajectory with no overlap with the training sequence. Results are averaged over 6 scenes.
We show an extensive study on depth and normal supervision and its effect on novel-view synthesis and depth metrics on challenging real-world scenes from the MuSHRoom and ScanNet++ datasets. Table 3 demonstrates that incorporating sensor depth supervision into 3DGS enhances both depth and RGB metrics compared to scenarios without geometric supervision. We qualitatively highlight the improvements in novel-view synthesis and depth quality, including the reduction of floaters and artifacts in Fig. 4.
Refer to caption (a) baseline Refer to caption (b) $\mathcal{L}_{\text{MSE}}$ Refer to caption (c) $\mathcal{L}_{1}$ Refer to caption (d) $\mathcal{L}_{\text{LogL1}}$ Refer to caption (e) $\mathcal{L}_{\hat{D}}$ (Ours) Refer to caption (f) iPhone GT
Figure 5: Qualitative comparison of depth losses: ScanNet++. We observe that the proposed gradient aware $\mathcal{L_{\hat{D}}}$ regularizer obtains the best qualitative results, mitigating uncertainties at edges from the raw iPhone depth captures. Zoom in to see the details.
@@
Splatfacto ($\mathcal{L}_{\text{rgb}}$[19, 45] 7.0 78.0 56.3 26.6 .870 12.9
@@
Table 4: Comparison of geometric supervision strategies. Geometric cues significantly improve reconstruction quality. We report mesh, novel view, and training time metrics (Nvidia 4090 GPU, 30k iterations) on the ’VR room’ sequence from MuSHRoom
Initialization Method # Init GS C-$L_{1}$downarrow NC $\uparrow$ F1$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
COLMAP SfM [41] $\sim$54.8K .0242 .8713 .9066 24.23 .8411
Sensor Depth 50 K .0242 .8725 .9074 24.16 .8389
Sensor Depth 100 K .0241 .8714 .9079 24.19 .8395
Sensor Depth 500 K .0239 .8730 .9091 24.20 .8399
Sensor Depth 1 M .0215 .8390 .9381 24.58 .8558
Sensor Depth 1.5 M .0238 .8730 .9171 24.31 .8491
Table 5: Comparison of SfM and sensor depth initialization. We compare Gaussian scene initialization using COLMAP [41] SfM points and those obtained from back-projecting training dataset sensor depth readings on the MuSHRoom dataset. We note that sensor depth initialization can improve overall reconstruction results compared to SfM initialization. We use 1M points in our experiments. Results are averaged over 6 scenes.
Method Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta<1.25$uparrow PSNR $\uparrow$ SSIM $\uparrow$
Splatfacto [45] .1481 .1345 .5122 .7602 23.41 .9111
Splatfacto + $\mathcal{L}_{\text{MSE}}$0551 .0180 .2104 .9647 23.84 .9140
Splatfacto + $\mathcal{L}_{\mathrm{1}}$0351 .0170 .1889 .9758 23.74 .9138
Splatfacto + $\mathcal{L}_{\mathrm{LogL1}}$0364 .0180 .1955 .9744 23.77 .9139
Splatfacto + $\mathcal{L}_{\hat{D}}$0285 .0162 .1790 .9782 23.61 .9122
Table 6: Quantitative comparison of depth losses: ScanNet++. We compare various depth supervision strategies and the resulting novel-view and mesh reconstruction results. See Fig. 5 for qualitative comparisons of the losses.
@@
Figure 6: Qualitative comparison of normal supervision. We observe that using a direct normal loss $\mathcal{L}_{\hat{N}}$ (a) results in non-smooth surfaces whereas using a normal smoothing prior $\mathcal{L}_{\text{smooth}}$ (b) significantly improves normal estimates according to (c) Omnidata  [7] predictions.
Method C-$L_{1}$downarrow NC $\uparrow$ F1$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
2DGS .0687 .8008 .6039 22.52 .8185
2DGS w/ Monodepth [3] .0795 .8558 .5446 21.85 .8009
2DGS w/ Sensor Depth .0275 .8830 .8886 23.02 .8250
Splatfacto .0652 .7727 .5835 24.44 .8486
Ours w/ Monodepth [3] .0477 .0860 .6963 24.48 .8509
Ours w/ Sensor Depth .0216 .8822 .9243 24.58 .8558
Table 7: 2DGS vs. DN-Splatter. We implement depth supervision to the recent 2DGS [16] method and compare with monocular and sensor depth regularization on MuSHRoom. We use the same loss and initialization strategy for fair comparison. Results are averaged over 6 scenes. NVS is reported within the train sequence.
5.3 Ablation studies
Proposed regularization strategy. We evaluate various design choices of our method in Table 4. The normal loss $\mathcal{L}_{\hat{N}}$ (Eq. 10) helps align Gaussians along the scene geometry improving the resulting mesh normal-completeness and F-scores. The depth loss $\mathcal{L}_{\hat{D}}$ (Eq. 4) significantly improves reconstruction quality and novel-view synthesis in ambiguous, textureless regions which are common in indoor scenes. The normal smoothing prior $\mathcal{L}_{\text{smooth}}$ (Eq. 11) further helps normal-completeness for Poisson meshing with a minimal impact on other metrics. Although the smoothing prior’s effect on quantitative metrics is minimal, its importance is evident in the qualitative renders illustrated in Fig. 6.
SfM vs. sensor depth initialization. We compare initialization strategies using sparse COLMAP [41, 40] SfM points and those obtained from back-projecting sensor depth readings. We note that using dense sensor depth data for initialization enhances both mesh and novel-view synthesis metrics, demonstrating the value of incorporating available sensor depth data.
2DGS vs. DN-Splatter. We evaluate our method against a variant of 2DGS [16] with the same depth regularization strategy in Table 7. Our comparison includes both monocular depth supervision with the Pearson correlation loss [52] and sensor depth supervision. We observe that 2DGS also faces difficulties with textureless regions, which are common in indoor datasets, leading to challenges with planar surfaces.
Depth supervision and losses. We assess the impact of the proposed depth loss Eq. 4 on ScanNet++, with qualitative renders shown in Fig. 5 and quantitative results in Table 6. We compare common depth losses: $\mathcal{L}_{\text{MSE}}$, $\mathcal{L}_{\mathrm{1}}$, $\mathcal{L}_{\mathrm{LogL1}}$ [14], and our edge-aware $\mathcal{L}_{\hat{D}}$. Results indicate that $\mathcal{L}_{\mathrm{1}}$ and $\mathcal{L}_{\mathrm{LogL1}}$ generally yield the best color metrics, with the logarithmic variant providing the smoothest reconstructions. Further details and comparisons with ground truth Faro lidar scans are available in the supplementary material.
Lastly, Table 8 evaluates our regularization strategy against common alternatives. We compare reconstruction with SotA monocular [3, 15] and multi-view [60] networks using the alignment strategy outlined in Section 4.1 using Eq. 5. We also examine the patch-based Pearson correlation loss [52] for relative depth supervision. While it performs better than naive monocular depth supervision, reconstruction quality is still inferior to using iPhone depths. Despite their low resolution and inaccuracies (mainly at object edges), sensor depths remain practical for real-world indoor scenes. Further research is needed to improve monocular depth supervision performance.
Method Acc. $\downarrow$ Comp. $\downarrow$ C-$L_{1}$downarrow NC $\uparrow$ F-$\uparrow$